# Enhancing Human-Robot Collaborative Assembly in Manufacturing Systems Using Large Language Models

### v1

@@ I Introduction
Further advancing the field of HRC in manufacturing systems by improving the interaction of human operators and robots presents significant challenges. Specifically, interaction with robots induces psychological stress and tension among operators due to language barrier [8]. In contemporary manufacturing systems, human operators require extensive pre-service training and complex code development to ensure accurate and safe production with the robots [9]. These difficulties highlight the need to develop human-robot communication systems that enable interaction between humans and robots without extensive robotics training (C1). Another challenge involves the need for greater flexibility and adaptability in human-robot interaction. As the number of interactions between humans and robots on the shop floor increases, there is a higher chance of encountering unexpected changes and errors. Thus, HRC needs to be more adaptable to real-time changes and errors during the manufacturing assembly process (C2). Additionally, the demand for adaptability and flexibility extends to the need for human-robot collaborative assembly applications to integrate advanced technologies in a human-centric design to enhance communication and usability (C3).
Large Language Models (LLMs) have recently been introduced in the AI community to enhance natural language understanding and generation capabilities. These can be extended to improve human-robot interaction in the manufacturing facility. Models such as OpenAI’s GPT-3 [10] and GPT-4 [11], have shown proficiency in processing, understanding, and communicating in natural language. The integration of LLMs facilitates natural language communication between humans and robots. Using voice interaction for this communication enhances collaboration and operator safety in dynamic workspaces. The main contributions of this work are: (1) the use of LLMs for interpreting natural language to allow operators to coordinate with robotic arms, (2) a framework for system integration of voice commands, robotic arms, and vision systems to facilitate HRC in an assembly task, enhancing operational flexibility, and (3) a dynamic adaptation approach to task variations, errors, and obstacles through human-robot communication.
@@ II-A Human-Robot Collaboration in Manufacturing
Recently, a wide range of methods has been developed in the field of HRC to improve the safety and efficiency of interactions between humans and robots. For instance, Mendes et al. [12] proposed an easy-to-perform HRC operating space that is more accessible to human operators. Mendes et al. integrate task-based programming and various interaction modes, including gestures and physical contact. Another work by Meziane et al. [13] developed a system for collision-free trajectories for robots in shared workspaces, utilizing a neural network for obstacle avoidance. Furthermore, the research conducted by Zanchettin et al. [14] shows that an algorithmic approach to HRCs can predict human activity patterns to control the robot more safely and effectively, resulting in a significant reduction in task time. Thus, several studies in the field of HRC have enhanced worker productivity and safety by combining the human experience with the automation capabilities of robots. Moreover, some approaches have utilized environmental data to enhance safety and efficiency. For example, Fernandez et al. [15] developed a dual-arm robotic system with multisensor capabilities for safe and efficient collaboration, integrating gesture recognition. Additionally, Wei et al. [16] gather information on human operators and the environment using RGB-D videos and predict human intentions based on a deep-learning method.
Research efforts have introduced innovative methods, such as gesture and physical contact for interaction, and neural networks for dynamic obstacle avoidance. However, there has been limited research focusing on human-robot collaborative assembly in manufacturing systems that effectively integrates interaction between humans and robots with natural language capabilities. We aim to integrate an LLM-based approach to facilitate communication into the human-robot collaborative assembly framework leveraging the strengths of both human flexibility and robot precision within the dynamic manufacturing environment.
@@ II-B Large Language Models for Robots
Recent research has explored the integration of LLMs for robotic applications to improve human-robot interaction and robot functionalities [17, 18, 19, 20, 21, 22]. This includes the conversion of high-level human instructions into tasks that robots execute. The research by Huang et al. [17] explored leveraging LLMs, such as GPT-3 and CODEX, to convert high-level human instructions into executable actions for robots. Similarly, Singh et al. [18] developed a method called ProgPrompt. This method generates Python code for task planning that accounts for environmental conditions and pre-programmed actions. Furthermore, Lin et al. [19] developed a framework for taking natural language instructions and then developing task and motion plans. Other works in robotics have utilized LLMs to maximize the robot’s capability to comprehend and interact with its environment to resolve errors. Mees et al [20], proposed an approach to combine natural language input with visual information based on a multi-camera system to enhance task completion ability. Zhang et al. [21] utilizes LLM to address multi-agent cooperation challenges, resolving issues related to decentralized control, raw sensory data processing, and efficient task execution in embodied environments. A framework called KnowNo [22] enhances robot autonomy and efficiency in ambiguous tasks using conformal prediction to guide robots in seeking human help when faced with uncertainty.
These research studies demonstrate how integrating LLMs into robotic applications improves human-robot interaction and robot capabilities. The methodologies involve converting human instructions into executable robot actions and accounting for environmental conditions. However, existing studies have not addressed common challenges found in the manufacturing environment, such as adapting natural language communication for different operators (e.g., adapting commands based on skills, shift time, or other metrics), matching natural language to structured task sequences found in a manufacturing environment, and managing the timing of commands (e.g., identifying when to request operator assistance in the process). We aim to bridge the gap by improving the integration of natural language for human-robot interaction for an assembly process through an LLM-based framework.
@@
In this framework, a human can initialize and provide a set of tasks $T=\{t_{1},t_{2},\ldots,t_{n}\}$ as the objectives to be achieved by the robot. The robot’s capabilities $C=\{c_{1},c_{2},\ldots,c_{l}\}$ represent the specific actions that are programmed to perform within an assembly process. If the task matches the robot’s capabilities, the robot should proceed to execute the specified task $t_{i}$. That task, $t_{i}$, is broken down into subtasks (${t_{i1},t_{i2},\ldots,t_{ik}}$) that outline the sequential steps for completion. The framework has to track the last completed subtask, denoted as $t_{i\text{c}}$, where $1\leq c\leq k$ indicates the completed process within ${t_{i1},t_{i2},\ldots,t_{ik}}$. When an error is resolved by a human operator, the robot resumes from $t_{i\text{c}+1}$, ensuring continuous task progression. To enable this human-robot collaborative assembly process, this framework is structured into two primary layers: the physical layer and the virtual layer, bridging the gap between human inputs and robot actions.
@@ III-A Physical Layer
The physical layer facilitates the interaction between the human operator and the robot within a shared physical space based on the information and data gathered from the virtual layer. The physical consists of three primary components: the human commands, the robot actions, and the sensor data. A human command is where an operator provides voice instructions to control robot actions in executing various tasks. The robot action involves the robot performing actions based on a predefined set of $T$. The sensor data is utilized to monitor environmental conditions. This data ensures the action adapts to changes in the workspace (e.g. position, orientation, or availability of parts).
When a predefined event or error $e_{i}\in E$ is detected, the robot utilizes the communication protocol to alert the human operator. This is achieved through the LLM module, which converts $e_{i}$ into a natural language message $M_{\text{e}_{i}}(t_{i})$, communicated via text-to-speech technology. Different instances of $e_{i}$ can lead to varied message $M_{\text{e}_{i}}(t_{i})$ depending on the task $t_{i}$ error. After receiving and understanding $M_{\text{e}_{i}}(t_{i})$, the human operator resolves the corresponding $e_{i}$ and commands the robot to resume $t_{i}$.
@@ III-B Virtual Layer
The virtual layer serves to facilitate communication in natural language. This layer stores the system’s functionality to enable interaction between human commands and robot actions. The virtual layer of the system consists of two main agents: a human agent and a robot agent.
@@ III-B1 Human Agent
We have integrated a human agent into our system to enhance the human-robot interaction based on the approach by Zhang et al. [23]. The goal of the human agent in our proposed framework is to process the human commands in voice instructions into a text that the robot agent can understand. This agent ensures that the human command intentions are transferred to the robot agent without the need for manual programming. The human agent contains an Automatic Speech Recognition (ASR) module responsible for processing the human command voice data, which converts this data into text that can be processed by the robot agent. The communication module in the human agent serves as a bridge, transmitting commands and information to the robot.
@@ III-B2 Robot Agent
  • LLM Module: A LLM is used to interpret human commands into a corresponding task $t_{i}\in T$. Human commands and functional capabilities $C$ including the detailed functional information of $C$ are put into the LLM. The LLM then analyzes the language and context of the human command. This process allows the system to match the command with the predefined functional capabilities $C$, ensuring the intended task is accurately executed. The LLM module also communicates any detected issues from the task control module into understandable natural language for the human operators, facilitating HRC for error resolution.
@@ III-C Human-Robot Collaborative Assembly Workflow
The overall workflow is depicted in a sequence diagram,  Figure 2, which visually represents the process flow within a human-robot collaborative assembly framework. The sequence diagram shows how voice commands from the human operator are processed and interpreted by the LLM Module and guides the robot actions.
The process begins with a human operator delivering a voice command, which is converted into a discrete set of tasks $T$ by the LLM module that the robot agent is programmed to understand and execute. The robot then requests sensor data to execute $t_{i}$. If the data provided by the sensor is valid, the robot proceeds to execute the assigned $t_{i}$. The sensor module determines data validity by comparing detected parameters to predefined criteria. A successful task execution results in a completion message $M_{\text{c}}(t_{i})$ to the human operator via the LLM Module.
When sensor data is invalid, or $t_{i}$ has any errors, the robot agent initiates the communication protocol to generate an error message $M_{\text{e}_{i}}(t_{i})$ via the LLM module, aiming to inform the human operator of the specific error and its occurrence within subtask $t_{i\text{c}+1}$ for efficient resolution. Following the error identification and correction by the human operator, a new command by the human operator is issued to the robot. The robot then resumes task execution at $t_{i}$, starting from the interrupted subtask $t_{i\text{c}+1}$, based on the new sensor data. This procedure is repeated until $t_{i}$ is successfully completed.
@@ IV Case Study
The physical setup is implemented using a Six Degrees of Freedom (6DOF) UFactory xArm with a UFactory 2-finger gripper as the end effector [24]. The base of the robotic arm remains fixed at the center, accompanied by a mat on one side, while the parts are assembled on the adjacent side. Two fixed cameras as visual sensors were placed in the surrounding area, one focused on the mat area for pre-assembly component placement, and the other camera on the assembling process. The conversion between camera coordinates and robot arm coordinates is accomplished through the coordinate transformation, as described in Section IV-C.
@@ IV-A LLM and ASR Module
The LLM module, enabled by OpenAI’s GPT-4.0 [11], is the decision-making component of the proposed framework. The LLM module converts instructions from the human operator into tasks $T$ executable by the robot, utilizing its capabilities $C$ through OpenAI’s function calling feature [27]. This module ensures that the robot actions match the human operator commands, allowing for adaptive and responsive task management within the manufacturing environment.
@@ IV-B Sensor Module: Vision System
We incorporated a vision system as our sensor module. During the collaborative assembly process, the vision system provides feedback on environmental data to the task control module. This enables precise object detection and identification of object orientation, facilitating task execution and error management. To enable object detection, Yolov5, a computer vision model, is utilized [28]. Custom Yolov5 models are trained using a dataset assembled for object detection. This dataset comprises images of individual parts (i.e. housing, wedge, spring, end cap), with manual bounding box annotations around each part in the images, and distinct classes for each part are created. The vision system utilizes this dataset with bounded boxes as a reference to identify the parts, as shown in Figure 4. If a part is detected, the coordinates of the top-left and bottom-right corners of the bounding box are retrievable. Once these coordinates are obtained, the x and y coordinates of these points are utilized to determine the object’s midpoint. This midpoint serves as the target point for the robotic arm to pick up the part. Following the successful object identification process, the valid sensor data is transferred to the task control module for guiding robot action. For parts that are not accurately detected, the sensor module transfers invalid data to the task control module to notify the human operator via LLM.
@@ IV-C Task Control Module: Assembly Task
The objective of the Task Control Module is to execute tasks as interpreted from human commands by LLM and to manage errors within a human-robot collaborative assembly system. Error handling is also a critical function of this module. The task control module verifies sensor data from the vision system. For instance, this module checks if a component is properly positioned for a pickup. If the sensor data is valid, the task control module continues with its operation. When the sensor data is invalid (e.g. missing component), the module transfers detailed error information to the LLM module to convert the information into natural language. This process activates the communication protocol, allowing the human operator to recognize and address the error.
To enable the assembly process, the task control module integrates various functionalities, including coordinate transformation and position and rotation control, based on the sensor data obtained from the vision system. These functions enable the execution of tasks interpreted by the LLM. The coordinate transformation is conducted to align the robotic arm ($A$) and camera ($C$) coordinate systems within the assembly process’s base frame ($B$) to ensure an accurate assembly task. The transformation between the robot’s base (${}_{B}^{A}p$) and camera’s (${}_{B}^{C}p$) positions is defined by
${}_{B}^{A}p=_{C}^{A}T\cdot_{B}^{C}p,$ (1)
where ${}_{C}^{A}T$ is the transformation matrix incorporating translation and rotation. The matrix includes a 3x3 rotation matrix ${}_{C}^{A}R$ and a translation vector $t$, translating coordinates from the camera to the robotic arm system. The effective calculation of positions $(x,y)$ relative to the base frame relies on this transformation for the robotic arm’s precise movements and interactions within the assembly setup. The position and rotation control function is then implemented to fine-tune both the positional alignment and the orientation of components for assembly tasks. To translate camera-detected object coordinates ($x_{p},y_{p}$) from pixel values to the world coordinate system, We used predefined offset values ($x_{o},y_{o}$), calculated as
$V=x_{o}\times x_{p},\quad H=y_{o}\times y_{p},$ (2)
where $V$ and $H$ adjust the arm’s trajectory based on the visual sensor input. Concurrently, the rotation control mechanism assigns specific trajectory angles to the robotic gripper, ensuring precise orientation alignment of parts for insertion. The cable shark assembly process is shown in Figure 5.
@@ IV-D Case Study Results
Based on these scenarios, the evaluation focused on assessing the system’s proficiency in interpreting and performing tasks derived from human commands, particularly considering the linguistic and human diversity. The variations in phrasing and terminology across different commands were cataloged, as presented in Table I, to validate the system’s adaptability to diverse linguistic inputs. The instruction type was classified into three distinct categories: specific, moderately specific, and least specific. Five unique variations of each instruction type were provided across three distinct tasks. To ensure the reliability of the evaluation, each language variation scenario presented in Table I is conducted with three repetitions.
@@
The results highlight the effectiveness of integrating LLM into the human-robot collaborative assembly framework. Figure 6 showcases the human-robot communication utilizing the vision system, outlined in Section IV-B, to ensure task completion in each scenario. In scenario 1, the robot detects overlapping components which signal to human operators for help. Upon resolution, the human operator prompts task continuation. Similarly in scenario 2, the robot identifies a misassembled wedge, notifying the human operator, and then the human commands the robot to continue after manually assembling the wedge. For Scenario 3, the robot flags a missing spring component, requesting human intervention, and the human instructs the robot to resume the assembly once the issue is fixed.
Based on the evaluation setup, we assessed the system’s capability to understand and execute commands with varied language expressions. For tasks with specific instructions, success rates were high, averaging approximately 98%. The rates decreased to 76% with moderately specific instructions and dropped further to 36% with least specific instructions. The success rates for each category of instruction type are detailed in Table II. This data suggests a positive correlation between instruction type and task execution success.

### v2

@@ I Introduction
Further advancing HRC in manufacturing systems by improving the interaction of human operators and robots presents significant challenges. Specifically, interaction with robots induces psychological stress and tension among operators due to language barrier [8]. In contemporary manufacturing systems, human operators require extensive pre-service training and complex code development to ensure accurate and safe production with the robots [9]. These difficulties highlight the need to develop human-robot communication systems that enable interaction between humans and robots without extensive robotics training (C1). Another challenge involves the need for greater flexibility and adaptability in human-robot interaction. As the number of interactions between humans and robots on the shop floor increases, there is a higher chance of encountering unexpected changes and errors. Thus, HRC needs to be more adaptable to changes and errors during the manufacturing assembly process (C2). Additionally, human-robot collaborative assembly applications must integrate advanced technologies with a human-centric design to improve communication and usability (C3).
Large Language Models (LLMs) have recently been introduced in the AI community to enhance natural language understanding and generation capabilities. These can be extended to improve human-robot interaction in the manufacturing facility. Models such as OpenAI’s GPT-3 [10] and GPT-4 [11], have shown proficiency in processing, understanding, and communicating in natural language. The integration of LLMs facilitates natural language communication between humans and robots. Using voice interaction for this communication enhances collaboration and operator safety in dynamic workspaces. The main contributions of this work are: (1) the use of LLMs for interpreting natural language to allow operators to coordinate with robotic arms, (2) a framework for system integration of voice commands, robotic arms, and vision systems to facilitate HRC in an assembly task, enhancing operational flexibility, and (3) an adaptation to task errors and obstacles through human-robot communication in a manufacturing environment.
@@ II-A Human-Robot Collaboration in Manufacturing
Recently, a wide range of methods has been developed in the field of HRC to improve the safety and efficiency of interactions between humans and robots. For instance, Fernandez et al. [12] developed a dual-arm robotic system with multisensor capabilities for safe and efficient collaboration, integrating gesture recognition. Additionally, Wei et al. [13] gather information on human operators and the environment using RGB-D videos and predict human intentions based on deep-learning methods.
Some approaches have utilized multi-modal strategies, including natural language, to improve communication for HRC in manufacturing. For example, Liu et al. [14] focus on integrating various modalities, including speech commands, hand motions, and body motions to improve HRC. While it addresses speech command recognition using deep learning models, the interaction is limited to basic speech command recognition and does not focus on context-aware communication. Additionally, Wang et al. [15, 16] employ a teaching-learning model that includes natural language instructions to predict human intentions and facilitate collaboration. While this model uses natural language for multimodal processing, it does not emphasize utilizing language variations in interaction.
Previous research have introduced methods using environmental data to enhance safety and efficiency and natural language to improve HRC in manufacturing. However, there is limited research on human-robot collaborative assembly that effectively integrates natural language capabilities for context-aware communication and handling language variations. We aim to integrate an LLM-based approach to improve communication between humans and robots. Our approach is an initial step towards enhancing human-robot interaction by combining existing technologies, such as computer vision and LLM, to leverage human flexibility and robot precision in manufacturing.
@@ II-B Large Language Models for Robots
Recent research has explored integrating LLMs in robotic applications to enhance human-robot interaction and robot functionalities [17, 18, 19, 20, 21, 22]. This includes converting high-level human instructions into tasks robots can execute. Huang et al. [17] used LLMs, such as GPT-3 and CODEX, to convert high-level human instructions into executable actions for robots. Similarly, Singh et al. [18] developed a method called ProgPrompt, which generates Python code for task planning that accounts for environmental conditions and pre-programmed actions. Furthermore, Lin et al. [19] developed a framework for taking natural language instructions and then developing task and motion plans. Other works in robotics have utilized LLMs to maximize the robot’s capability to comprehend and interact with its environment to resolve errors. Mees et al [20] combined natural language input with visual information to enhance task completion ability. Zhang et al. [21] utilizes LLM to address multi-agent cooperation challenges, resolving issues related to decentralized control, raw sensory data processing, and efficient task execution in embodied environments. KnowNo [22] improved robot autonomy and efficiency by guiding robots in seeking human help when faced with uncertainty.
These studies show how integrating LLMs into robotic applications improves human-robot interaction and robot capabilities. However, existing studies have not addressed common challenges found in the manufacturing environment, such as matching natural language to structured task sequences and managing the timing of commands (e.g., identifying when to request operator assistance in the process). We aim to bridge the gap by improving the integration of natural language for human-robot interaction for an assembly process through an LLM-based framework.
@@
In this framework, a human can initialize and provide a set of tasks $T=\{t_{1},t_{2},\ldots,t_{n}\}$ as the objectives to be achieved by the robot. While $T$ is predefined, the system can be extended to integrate new tasks based on human input. The robot’s capabilities $C=\{c_{1},c_{2},\ldots,c_{l}\}$ represent the specific actions programmed to perform within an assembly process. If the task matches the robot’s capabilities, the robot should execute the specified task $t_{i}$. The task, $t_{i}$, is broken down into subtasks (${t_{i1},t_{i2},\ldots,t_{ik}}$) that outline the sequential steps for completion. The framework has to track the last completed subtask, denoted as $t_{i\text{c}}$, where $1\leq c\leq k$ indicates the completed process within ${t_{i1},t_{i2},\ldots,t_{ik}}$. When an error is resolved by a human operator, the robot resumes from $t_{i\text{c}+1}$, ensuring continuous task progression. To enable this human-robot collaborative assembly process, this framework is structured into two primary layers: the physical layer and the virtual layer, bridging the gap between human inputs and robot actions.
@@ III-A Physical Layer
The physical layer facilitates human-robot interaction within a shared space, based on data from the virtual layer. The physical consists of three primary components: human commands, robot actions, and sensor data. Human commands involve the operator providing voice instructions to control robot actions. Robot actions are based on a predefined set of $T$. Sensor data is utilized to monitor environmental conditions. This data ensures the action adapts to changes in the workspace (e.g. position, orientation, or availability of parts).
When a predefined event or error $e_{i}\in E$ is detected, the robot alerts the human operator using a communication protocol. The LLM module converts $e_{i}$ into a natural language message $M_{\text{e}_{i}}(t_{i})$, delivered via text-to-speech technology. Different instances of $e_{i}$ can lead to varied message $M_{\text{e}_{i}}(t_{i})$ depending on the task $t_{i}$ error. After receiving and understanding $M_{\text{e}_{i}}(t_{i})$, the human operator resolves the corresponding $e_{i}$ and commands the robot to resume $t_{i}$.
@@ III-B Virtual Layer
The virtual layer serves to facilitate communication. This layer stores the system’s functionality to enable interaction between human commands and robot actions. The virtual layer of the system consists of two main agents: a human agent and a robot agent.
@@ III-B1 Human Agent
We have integrated a human agent, often referred to as a human command interpreter, to enhance the human-robot interaction based on the approach by Zhang et al. [23]. The goal of the human agent is to process the human commands in voice instructions into a text that the robot agent can understand. This agent ensures that the human intentions are transferred to the robot without manual programming. The human agent contains an Automatic Speech Recognition (ASR) module responsible for processing the human command voice data, which converts this data into text that the robot agent can process. The communication module in the human agent serves as a bridge, transmitting commands and information to the robot.
@@ III-B2 Robot Agent
  • LLM Module: An LLM interprets human commands into tasks $t_{i}\in T$ using initialization prompts and detailed functional information. By processing commands based on context, such as the last completed subtask and any encountered errors, the LLM can automatically detect and suggest the next task. This process allows the system to match the command with the predefined functional capabilities $C$, ensuring the intended task is executed. The LLM module also communicates any detected issues from the task control module into understandable natural language for the human operators, facilitating HRC for error resolution. Using an LLM to process static error messages can generate contextual information, helping human operators understand and resolve issues.
@@ III-C Human-Robot Collaborative Assembly Workflow
The overall workflow is shown in a sequence diagram in  Figure 2, illustrating the human-robot collaborative assembly process. The diagram depicts how voice commands from the human operator are processed by the LLM module to guide robot actions.
The process starts with the operator giving a voice command, converted by the LLM module into a discrete set of tasks $T$ for the robot. The robot then requests sensor data to execute $t_{i}$. If the data is valid, the robot proceeds to execute the assigned $t_{i}$. The sensor module determines data validity by comparing detected parameters to predefined criteria. Successful execution results in a completion message $M_{\text{c}}(t_{i})$ to the operator via the LLM Module.
If the data is invalid, or $t_{i}$ has any errors, the robot generates an error message $M_{\text{e}_{i}}(t_{i})$ via the LLM module, aiming to inform the human operator of the specific error and its occurrence within subtask $t_{i\text{c}+1}$ for efficient resolution. Following the error identification and correction by the human operator, a new command by the human operator is issued to the robot. The robot then resumes task execution at $t_{i}$, starting from the interrupted subtask $t_{i\text{c}+1}$, based on the new sensor data. This procedure is repeated until $t_{i}$ is completed.
@@ IV Case Study
The physical setup uses a Six Degrees of Freedom (6DOF) UFactory xArm with a UFactory 2-finger gripper as the end effector [24]. The base of the robotic arm remains fixed at the center, accompanied by a mat on one side, while the parts are assembled on the adjacent side. Two Intel RealSense Depth Camera D435 cameras were used as visual sensors in the surrounding area, one focused on the mat area for pre-assembly component placement, and the other camera on the assembling process. The conversion between camera and robot arm coordinates is accomplished through the coordinate transformation. We used a fixed z-axis value for the robotic arm to pick up camera-identified parts. In future work, we aim to make this z-axis value adaptable.
@@ IV-A LLM and ASR Module
The LLM module leverages OpenAI’s pre-trained GPT-4.0 [11]. The LLM module converts instructions from the human operator into tasks $T$ executable by the robot, utilizing its capabilities $C$ through OpenAI’s function calling feature [27]. This module ensures that the robot actions match the human operator commands, allowing for adaptive and responsive task management within the manufacturing environment.
@@ IV-B Sensor Module: Vision System
We incorporated a vision system as our sensor module. During the collaborative assembly process, the vision system provides feedback on environmental data to the task control module. This enables precise object detection and identification of object orientation, facilitating task execution and error management. To enable object detection, Yolov5, a computer vision model, is utilized [28]. Custom Yolov5 models are trained using a dataset of images of individual parts (i.e. housing, wedge, spring, end cap), with manual bounding box annotations around each part in the images. The vision system these annotations to identify the parts, as shown in Figure 4. If a part is detected, the coordinates of the top-left and bottom-right corners of the bounding box are retrievable. Once these coordinates are obtained, the x and y coordinates of these points are utilized to determine the object’s midpoint. This midpoint serves as the target point for the robotic arm to pick up the part. Following the successful identification, the valid sensor data is transferred to the task control module for guiding robot action. For inaccurately detected parts, the sensor module transfers invalid data to the task control module to notify the human operator via LLM.
@@ IV-C Task Control Module: Assembly Task
The Task Control Module executes tasks from human commands interpreted by the LLM and manages errors in a human-robot collaborative assembly system. The task control module verifies sensor data from the vision system. For instance, if the data is valid, the task control module proceeds; if not, the task control module sends error details (e.g. missing component) to the LLM module to convert the information into natural language. This process activates the communication protocol, allowing the human operator to recognize and address the error. To enable the assembly process, the task control module integrates various functionalities, including coordinate transformation and position and rotation control, based on the sensor data obtained from the vision system. These functions enable the execution of tasks interpreted by the LLM. The orientation of detected objects is determined from the bounding box dimensions, assuming the object’s angle is perpendicular or parallel to the x-axis of the camera vision. The cable shark assembly process is shown in Figure 5.
@@ IV-D Case Study Results
Based on these scenarios, the evaluation assessed the system’s proficiency in interpreting and performing tasks from human commands, considering the linguistic and human diversity. Variations in phrasing and terminology across different commands were cataloged, as presented in Table I, to validate the system’s adaptability to diverse linguistic inputs. Instructions were classified into three distinct categories: specific, moderately specific, and least specific. Five unique variations of each instruction type were provided across three distinct tasks. To ensure reliability, each language variation scenario presented in Table I is conducted with three repetitions.
@@
The results highlight the effectiveness of integrating LLM into the human-robot collaborative assembly framework. Figure 6 showcases the human-robot communication using the vision system, outlined in Section IV-B, to ensure task completion. In scenario 1, the robot detects overlapping components which signal to human operators for help. Upon resolution, the human operator prompts task continuation. Similarly in scenario 2, the robot identifies a misassembled wedge, notifying the human operator, and then the human commands the robot to continue after manually assembling the wedge. For Scenario 3, the robot flags a missing spring component, requesting human intervention, and the human instructs the robot to resume the assembly once the issue is fixed.
The evaluation assessed the system’s capability to understand and execute commands with varied language expressions. Success rate refers to the percentage of the robot correctly understanding and executing the given instruction without requiring additional clarification. For tasks with specific instructions, success rates were high, averaging approximately 98%. The rates decreased to 76% with moderately specific instructions and dropped further to 36% with least specific instructions. The success rates are detailed in Table II. This data suggests a positive correlation between instruction type and task execution success.